OpenStudy (anonymous):

y varies directly with x. When y = 20, x = 50. Find x when y = 36. Choose one answer. a. 75 b. 80 c. 90 d. 105

OpenStudy (anonymous):

what do you get?

OpenStudy (saifoo.khan):

90 is your answer.

myininaya (myininaya):

y=kx 20=k(50) 20/50=k 2/5=k k=2/5 y=2x/5 36=2x/5 5/2 * 36=x 5*36/2=x 5* 18=x 90=x

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

thanks guys! thats what i was thinking!

myininaya (myininaya):

when you see varies think multiplication when you see inversely think division

OpenStudy (saifoo.khan):

uw.

myininaya (myininaya):

never forget your constant

OpenStudy (anonymous):

you can always forget the constant! when you see "varies directly" just look at how the one is varying, and vary the other the same. Here, y goes from 20 to 36, getting multiplied by 1.8 (36/20). So to vary x, we multiply it by 1.8. 50*1.8 = 90.

myininaya (myininaya):

no you need the constant to find what the constant is lol